OpenStudy (anonymous):

find horizontal asymptote of (x-2)/(x^+1) +2

OpenStudy (anonymous):

wouldn't it be just zero?

OpenStudy (anonymous):

sorry I made a mistake! the bottom is \[x ^{2}-1\]

OpenStudy (anonymous):

does the +2 at the end change anything?

OpenStudy (jdoe0001):

\(\bf \cfrac{x-2}{(x^2-1)+2} \ \ ?\)

OpenStudy (anonymous):

no the +2 is outside

OpenStudy (anonymous):

so its just 2?

OpenStudy (jdoe0001):

\(\bf \cfrac{x-2}{(x^2-1)}+2\implies \cfrac{x-2+2x^2-2}{x^2-1} \implies \cfrac{3x^2-4}{x^2-1}\)

OpenStudy (jdoe0001):

so the horizontal asymptote for a rational whose num/den are same degree is the leading coefficients of n/d

OpenStudy (jdoe0001):

the +2 will make the distinction that you'd need to consolidate first, into one rational

OpenStudy (anonymous):

wait how did that top become that?

terenzreignz (terenzreignz):

Whoops... jdoe shouldn't have added the x to the \(\large 2x^2\) Top should be as follows: \[\Large \frac{\color{green}{2x^2 + x-4}}{x^2 -1}\]

OpenStudy (anonymous):

why are you adding 2x^2

OpenStudy (anonymous):

or how did that top change?

OpenStudy (anonymous):

from x-2???

OpenStudy (ybarrap):

Is this the problem? $$\tt \begin{align*} \\ \frac{x-2}{x^2 -1}\\ \end{align*}$$

OpenStudy (anonymous):

yeah and +2

OpenStudy (ybarrap):

$$\tt \begin{align*} \\ \text{Horizontal Asymptote:}~ lim_{x\to\infty}\frac{x-2}{x^2-1}=0\\ \end{align*}$$

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

its \[[(x-2)/(x ^{2}-1)] +2\]

OpenStudy (anonymous):

\[\frac{ x-2 }{x^2+1}+2\]

OpenStudy (anonymous):

sorry that's the correct one

OpenStudy (ybarrap):

np $$\tt \dfrac{ x-2 }{x^2+1}+2\\\\ \text{Horizontal Asymptote:}~~ lim_{x\to\infty}\dfrac{ x-2 }{x^2+1}+2=2\\ $$

OpenStudy (anonymous):

horizontal asymptote is 2?

OpenStudy (anonymous):

why?

OpenStudy (ybarrap):

Yes. The asymptote is where the function will "settle" as x gets larger. In the long-run, the first part of your equation converges (or "settles") to 0, while as x gets larger, the second part of your equation, the "2", which is of course a constant, keeps the entire function at 2.

OpenStudy (ybarrap):

The first part of your equation has x with degree 1 on the numerator and this is the dominant term as x gets larger. In the denominator, the x has a degree 2. As x gets largerer, x^2 in the denominator will become grow faster than "x" in the numerator. So, when x gets big enough, this whole fractional part looks like x/x^2 or 1/x. So 1/x tends to zero as x gets larger, leaving the constant "2" to dominate and thus creating the horizontal asymptote. The whole thing tends to 2 as x gets larger, never quite getting there.

OpenStudy (anonymous):

ok thanks!